hi,
could you please give me the equations/solutions for these questions,
thanks,
1. the sum of two number is 20. what is the least possible sum of their squares?
2. what number exceeds its square by the greatest amount?
i will be waiting.
thank you

Answer by Alan3354(69443)

1. the sum of two number is 20. what is the least possible sum of their squares?
2. what number exceeds its square by the greatest amount?
--------------
1. S = x^2 + (20-x)^2
S = x^2 + 400 -40x +x^2
S = 2x^2 - 40x + 400
Set 1st derivative = 0
4x - 40 = 0
x = 10
10^2 + 10^2 = 200
--------------
2. D = x-x^2
Set 1st derivative = 0
1-2x = 0
x = 1/2